# What Makes Math Word Problems Challenging for LLMs?

### v1

@@ C.1 Preprocessing
    Scaling: We fit a scikit-learn’s StandardScaler onto the train split to normalize the mean and standard deviation of all features. We then apply the same scaler on the test split features.
@@
    Balancing: As many LLM solution split either contain too few always-correct or always-incorrect question samples, we use imblearn’s RandomOverSampling tool to balance the proportions of the two classes in each run.
@@ C.2 Hyperparameter Search
For each classfier model and LLM solution split pair, we conduct Bayesian optimization on the ranges of key hyperparameters for each classifier. As the objective function, we maximize the Macro-F1 score on a 15% held-out set of GSM8k’s test split to prevent the models from over-fitting onto the train samples.
@@ D.1 Class-Wise Classification Review
Table 8 reports the class-wise F1-scores (for questions that are always or never solved correctly) for each classifier across LLM question splits. Though there may be notable class imbalance among the two classes across splits, all classifiers were trained with proportional oversampling. We see that for relatively smaller pretrained models, i.e., Llama2-13B and Mistral-7B, F1 scores for always-incorrect questions are significantly higher than their counterpart. Thus, the scores indicate that for smaller models, questions answered incorrectly are more predictable. For larger models like Llama2-70B, this difference is lower, with the always-correct questions being somewhat more predictable. For the fine-tuned MetaMath-13B model, the small number of questions that are never answered correctly, fail to provide a generalizable sample for predicting on unseen data.
@@ D.2 Feature Impact
We continue our discussion of feature importance (from Section 4.2.1) by identifying pivot points for key features, about which the corresponding success rates for questions show a significant difference in mean value. We perform Student’s $t$-tests on equally spaced threshold along each feature and report the threshold which show the highest variation in mean success rate in Table 7.
We see that across most LLMs, a significant rise in mean success rate is observed as the question contains on average, more than $6$-$7$ sentences, $90$-$115$ words, or $113$-$132$ tokens. We get a better idea of the kind of questions Mistral-7B gets wrong more often than other models, as its threshold ($19.71$ words) for number of words in the question body is substantially lower than the average. Across features involving number of math operations, the threshold for the fine tuned MetaMath-13B model is either at par or higher than other models.

### v2

@@ C.1 Preprocessing
    Scaling: We fit scikit-learn’s StandardScaler onto the train split to normalize the mean and standard deviation of all features. We then apply the same scaler on the test split features.
@@
    Balancing: As many LLM solution splits either contain too few always-correct or always-incorrect question samples, we use imblearn’s RandomOverSampling tool to balance the proportions of the two classes in each run.
@@ C.2 Hyperparameter Search
For each classifier model and LLM solution split pair, we conduct Bayesian optimization on the ranges of key hyperparameters for each classifier. As the objective function, we maximize the Macro-F1 score on a 15% held-out set of GSM8K’s test split to prevent the models from over-fitting onto the train samples.
C.3 Fine-tuning RoBERTa Classifier
We use HuggingFace’s Trainer module to tune a pre-trained RoBERTa-base classifier on the same target data as the statistical classifiers in each setting. The corresponding input text for training and evaluation was built by concatenating the GSM8K question and gold-solution text for each sample, i.e., "<question> Question Body </question> <solution> Solution Body </solution>". The model is trained for $3$ epochs with a peak learning rate of $52e-$ 5 and a warmup ratio of $0.1$. On a Tesla P100, each training run took approximately 10-15 minutes.
@@ D.1 Class-wise Classification Review
Table 8 reports the class-wise F1-scores (for questions that are always or never solved correctly) for each classifier across LLM question splits. Though there may be notable class imbalances among the two classes across splits, all classifiers were trained with proportional oversampling. We see that for relatively smaller pretrained models, i.e., Llama2-13B and Mistral-7B, F1 scores for always-incorrect questions are significantly higher than their counterparts. Thus, the scores indicate that for smaller models, questions answered incorrectly are more predictable. For larger models like Llama2-70B, this difference is lower, with the always-correct questions being somewhat more predictable. For the fine-tuned MetaMath-13B model, the small number of questions that are never answered correctly, fail to provide a generalizable sample for predicting on unseen data.
@@ D.2 Feature Impact
We continue our discussion of feature importance (from Section 4.2.1) by identifying pivot points for key features about which the corresponding success rates for questions show a significant difference in mean values. We perform Student’s $t$-tests on equally spaced thresholds along each feature and report the thresholds which show the highest variation in mean success rates in Table 7.
We see that across most LLMs, a significant rise in the mean success rate is observed as the question contains, on average, more than $6$-$7$ sentences, $90$-$115$ words, or $113$-$132$ tokens. We get a better idea of the kind of questions Mistral-7B gets wrong more often than other models, as its threshold ($19.71$ words) for the number of words in the question body is substantially lower than the average. Across features involving number of math operations, the threshold for the fine-tuned MetaMath-13B model is either on a par or higher than other models.